# AgentTrace: Causal Graph Tracing for Root Cause Analysis in Deployed Multi-Agent Systems

### 16 Mar 2026 v1

@@ 1 Introduction
While our benchmark scenarios are synthetically constructed, they are designed to capture common failure patterns observed in deployed multi-agent systems, such as early planning errors cascading through execution. Our evaluation across 550 failure scenarios in 10 domains shows that AgentTrace achieves high localization accuracy with sub-second latency, significantly outperforming both heuristic and LLM-based baselines.
@@ 3.4 Node Ranking Algorithm
$F_{i}(v)=\frac{1}{|G_{i}|}\sum_{j\in G_{i}}\hat{f}_{j}(v)$ (2)
where $G_{i}$ is the set of features in group $i$ and $\hat{f}_{j}(v)\in[0,1]$ is the min-max normalized feature value (see Appendix B).
where each $F_{i}(v)$ aggregates normalized features for group $i$, and the weights $w_{i}$ are determined via grid search on a held-out validation set of 50 scenarios (details in Appendix B). The feature groups and their learned weights are:
Position Features ($w_{p}=0.70$).
Position-based features capture where the node occurs in the execution:
  • Normalized position: $\text{pos}(v)/|V|$
  • Distance to error: Hop count from $v$ to $v_{\text{error}}$
  • Depth in trace: Steps from the initial node
Structure Features ($w_{s}=0.20$).
Graph topology features indicating node importance:
  • Out-degree: Number of downstream nodes affected
  • Betweenness: Fraction of paths passing through $v$
  • Fanout ratio: Proportion of total nodes reachable from $v$
Content Features ($w_{c}=0.05$).
Semantic indicators of problematic content:
  • Error keywords: Presence of terms like “error”, “failed”
  • Uncertainty markers: Terms like “maybe”, “not sure”
  • Content length: Unusually short or long outputs
Flow Features ($w_{f}=0.03$).
Agent interaction patterns:
  • Agent switch: Whether node involves cross-agent communication
  • Role criticality: Importance of the agent’s role
Confidence Features ($w_{e}=0.02$).
Model-reported confidence when available:
  • Stated confidence: Explicit confidence scores
  • Hedging language: Implicit uncertainty indicators
@@ 4.1 Scenario Generation
We created a benchmark of 550 synthetic multi-agent failure scenarios across 10 domains (Table 1). Each scenario includes:
  • A complete execution trace with 8-15 agent actions
  • A systematically injected bug at a specific node
  • Ground-truth annotation of the root cause node
Table 1: Benchmark domains and scenario counts
Domain Scenarios
Software Development 52
Customer Service 51
Research Analysis 51
Planning & Scheduling 46
Financial Trading 50
Healthcare Coordination 60
Legal Document Analysis 60
Educational Tutoring 60
Financial Advisory 60
DevOps Automation 60
Total 550
4.2 Bug Types
We inject five categories of bugs:
  • Logic errors: Incorrect conditional decisions
  • Communication failures: Message misinterpretation
  • Data corruption: Incorrect data transformation
  • Missing validation: Skipped verification steps
  • Role confusion: Agent acting outside its expertise
4.3 Ground Truth Annotation
Each scenario’s ground truth was established through: (1) controlled bug injection at a known step, (2) verification that removing the bug prevents the error, and (3) confirmation that the bug’s effects propagate to the error node.
@@ 5.2 Baselines
  • LLM Analysis: GPT-4 with full trace and prompt engineering
@@ 5.3 Main Results
Table 2 shows that AgentTrace significantly outperforms all baselines.
Table 2: Root cause localization accuracy on 550-scenario benchmark. 95% confidence intervals from bootstrap resampling.
Method Hit@1 Hit@3 MRR
Random 9.1% 27.3% 0.18
First Node 3.6% 10.9% 0.07
Last Node 12.7% 38.2% 0.25
LLM Analysis 68.5% 81.4% 0.74
AgentTrace 94.9% 98.4% 0.97
[92.9, 96.7] [96.9, 99.4]
Statistical significance testing using McNemar’s test confirms that AgentTrace significantly outperforms all baselines ($p<0.001$). The effect size (Cohen’s h) compared to LLM Analysis is 0.77, indicating a large practical difference.
@@ 5.4 Ablation Study
Table 3 shows Hit@1 when using each feature group in isolation. Position features alone achieve 87.3% accuracy, confirming that bug location patterns are highly predictive in our benchmark.
Table 3: Ablation study: Hit@1 with each feature group alone
Feature Group Hit@1
Position only 87.3%
Structure only 34.5%
Content only 28.7%
Flow only 15.2%
Confidence only 12.1%
All features 94.9%
5.5 Performance by Domain
Table 4 shows consistent performance across domain groups, with technical domains (Software, DevOps) slightly outperforming others.
Table 4: Performance by domain group
Domain Group Hit@1 Scenarios
Technical (Dev, DevOps) 96.4% 112
Business (Trading, Finance) 93.6% 110
Service (Customer, Healthcare) 93.7% 111
Knowledge (Research, Legal, Education) 96.5% 171
Planning 91.3% 46
@@ 7 Conclusion
Our benchmark generation code, evaluation scripts, and all experimental results are available at https://github.com/zwang000/AgentTrace. The benchmark includes complete execution traces, ground truth annotations, and baseline implementations.
@@ Appendix A Benchmark Generation Details
Bug Injection.
We employ a five-step injection process: (1) generate a correct execution trace from domain templates; (2) select a bug location (early steps 2–3: 60%, middle steps 4–6: 30%, late steps 7+: 10%); (3) assign a bug type per the distribution in Table 6; (4) inject erroneous content (e.g., flipped comparison operators for logic errors, message truncation for communication failures, swapped variable names for data corruption); (5) propagate cascading effects to downstream steps. To prevent methods from exploiting explicit bug markers, we also produce a blind version with anonymized IDs and separated ground truth.
Table 6: Bug type distribution in benchmark
Bug Type Count Percentage
Logic Error 165 30.0%
Communication Failure 110 20.0%
Data Corruption 110 20.0%
Missing Validation 88 16.0%
Role Confusion 77 14.0%
Total 550 100%
Causal Graph Statistics.
The constructed causal graphs have on average 10.8 nodes ($\sigma{=}2.1$) and 14.2 edges ($\sigma{=}3.5$), with average degree 2.6, diameter 6.3, and mean path length 3.2. Edges are predominantly sequential (61.3%), followed by communication (26.8%) and data dependency (11.9%).
Appendix B Feature Extraction and Weight Learning
Table 7 lists all features used in the ranking algorithm. All features are normalized to $[0,1]$ via min-max normalization within each scenario:
$\hat{f}_{i}(v)=\frac{f_{i}(v)-\min_{u\in V}f_{i}(u)}{\max_{u\in V}f_{i}(u)-\min_{u\in V}f_{i}(u)+\epsilon}$ (3)
where $\epsilon=10^{-8}$ prevents division by zero.
Table 7: Complete feature list with computation methods
Group Feature Formula Range
Position Normalized Position $\text{pos}(v)/|V|$ $[0,1]$
Depth Ratio $\text{depth}(v)/\max_{u}\text{depth}(u)$ $[0,1]$
Reverse Position $1-\text{pos}(v)/|V|$ $[0,1]$
Structure Out-degree $|\{u:(v,u)\in E\}|/\max_{w}|\{u:(w,u)\in E\}|$ $[0,1]$
In-degree $|\{u:(u,v)\in E\}|/\max_{w}|\{u:(u,w)\in E\}|$ $[0,1]$
Reachability $|\text{descendants}(v)|/|V|$ $[0,1]$
Content Error Keywords $\mathbb{1}[\text{``error''}\in\text{content}(v)]$ $\{0,1\}$
Uncertainty $\mathbb{1}[\text{``maybe''}\in\text{content}(v)]$ $\{0,1\}$
Length Anomaly $|\text{len}-\mu|/(3\sigma)$ (clipped) $[0,1]$
Keyword Density $\text{count}(\text{keywords})/\text{len}$ $[0,1]$
Flow Agent Switch $\mathbb{1}[\text{agent}(v)\neq\text{agent}(\text{prev}(v))]$ $\{0,1\}$
Role Criticality $\text{role\_weight}(\text{agent}(v))$ $[0,1]$
Communication $\mathbb{1}[\text{type}(v)=\text{``message''}]$ $\{0,1\}$
Confidence Stated Confidence $\text{conf}(v)$ if available, else 0.5 $[0,1]$
Hedging Score $\text{count}(\text{hedge\_words})/10$ (clipped) $[0,1]$
Weight Learning.
@@ Appendix C Statistical Analysis Details
Table 8 provides the complete statistical comparisons.
Table 8: Complete statistical comparison of AgentTrace vs. baselines
Baseline Hit@1 $n_{01}$ $n_{10}$ $\chi^{2}$ $p$-value Cohen’s h
Random 9.1% 478 6 459.2 $10^{-10}$ 2.41
First Node 3.6% 503 2 497.0 $10^{-10}$ 2.72
Last Node 12.7% 460 8 436.5 $10^{-10}$ 2.29
LLM Analysis 68.5% 173 28 102.5 $10^{-10}$ 0.77
Appendix D Detailed Experimental Results
Table 9 provides per-domain performance. Tables 1012 stratify results by bug type, trace length, and bug position.
@@ Appendix D Detailed Experimental Results
Table 10: Performance by bug type
Bug Type N Hit@1 MRR
Logic Error 165 95.8% 0.98
Communication Failure 110 93.6% 0.96
Data Corruption 110 94.5% 0.97
Missing Validation 88 94.3% 0.96
Role Confusion 77 94.8% 0.97
Table 11: Performance by trace length
Trace Length N Hit@1 MRR
8–9 steps 132 95.5% 0.97
10–11 steps 198 95.0% 0.97
12–13 steps 154 94.2% 0.96
14–15 steps 66 93.9% 0.96
Table 12: Performance by bug position in trace
Bug Position N Hit@1 MRR
Early (2–3) 330 96.1% 0.98
Middle (4–6) 165 93.3% 0.96
Late (7+) 55 92.7% 0.95
Appendix E Ablation Study Details
Table 13 shows Hit@1 for all feature group combinations. Position features alone reach 87.3%, while each additional group provides incremental gains. Table 14 shows that the optimal position weight $w_{p}=0.70$ balances the strong predictive power of position features with discriminative information from other groups.
Table 13: Ablation study: Performance with different feature combinations
Feature Combination Hit@1 $\Delta$ vs Full
All features (P+S+C+F+E) 94.9%
Position only (P) 87.3% -7.6%
Structure only (S) 34.5% -60.4%
Content only (C) 28.7% -66.2%
Flow only (F) 15.2% -79.7%
Confidence only (E) 12.1% -82.8%
P + S 92.4% -2.5%
P + C 90.9% -4.0%
P + F 89.5% -5.4%
P + E 88.7% -6.2%
S + C 45.6% -49.3%
S + F 38.2% -56.7%
P + S + C 93.8% -1.1%
P + S + F 93.3% -1.6%
P + S + C + F 94.5% -0.4%
Table 14: Sensitivity of Hit@1 to position weight $w_{p}$
$w_{p}$ 0.50 0.60 0.70 0.80 0.90
Hit@1 91.3% 93.5% 94.9% 93.8% 91.1%
Appendix F LLM Baseline Details
We use GPT-4 (gpt-4-0613) with temperature 0.0, max tokens 1024, top-$p$ 1.0, and zero frequency/presence penalties. The prompt instructs the model to analyze the full execution trace, identify causal relationships, and output only the step number of the root cause:
@@ 4. Consider: logic errors, miscommunication,
Error Type Count Percentage
Selected error node instead of root cause 82 47.4%
Off-by-one (adjacent to correct node) 45 26.0%
Selected intermediate propagation step 31 17.9%
Completely incorrect 15 8.7%
Total Errors 173 100%
Appendix G Runtime Performance Details
Tables 17 and 17 break down runtime by component and show scaling with trace length. Runtime scales approximately linearly: $T\approx 12.5n+5$ ms, where $n$ is the number of steps. All experiments were conducted on an AMD Ryzen 9 5950X (16-core, 3.4 GHz), 128 GB DDR4 RAM, NVMe SSD, Ubuntu 24.04 LTS, Python 3.13.
@@ Appendix G Runtime Performance Details
Appendix H Example Scenarios
Example 1: Software Development.
Task: Implement average calculation.
Step 1 [Planner]: Analyze requirements
  -> "Sum all numbers, count elements, divide"
Step 2 [Coder]: Write implementation
  -> "def average(nums): ..."
Step 3 [Coder]: Add edge case handling  [BUG]
  -> "if nums = []:  # BUG: = instead of =="
Step 4 [Reviewer]: Review code
  -> "Code looks correct."
Step 5 [Executor]: Run tests  [ERROR]
  -> "SyntaxError: invalid syntax at line 2"
Root cause: Step 3 (logic error). AgentTrace correctly identifies Step 3; LLM selects Step 5 (error node).
Example 2: Customer Service.
Task: Handle delayed shipment complaint.
Step 1 [Router]: Classify request -> "Shipping Issue"
Step 2 [Specialist]: Look up order -> "In Transit"
Step 3 [Specialist]: Misread delivery date  [BUG]
  -> "Expected: March 15 (actually March 5)"
Step 4 [Resolver]: Compose response
  -> "On track for March 15..."
Step 5 [Logger]: Send response  [ERROR]
  -> "Customer: ’Was supposed to arrive March 5!’"
Root cause: Step 3 (data corruption). Both AgentTrace and LLM correctly identify Step 3.
Example 3: Research Analysis.
Task: Analyze transformer architecture publications.
Step 1 [Searcher]: Query databases -> "150 papers"
Step 2 [Searcher]: Filter -> "25 relevant papers"
Step 3 [Analyzer]: Extract findings  [BUG]
  -> "Attention being replaced by MLPs"
     (actually: attention being enhanced)
Step 4 [Synthesizer]: "Field moving away from
     attention-based models..."
Step 5 [Writer]: "Transformers becoming obsolete"
Step 6 [Writer]: Finalize report  [ERROR]
  -> "Contradicts recent SOTA results"
Root cause: Step 3 (communication failure). AgentTrace correctly identifies Step 3; LLM selects Step 4 (propagation step).
Appendix I Failure Case Analysis
Among the 28 scenarios (5.1%) where AgentTrace failed to rank the correct root cause first, Table 18 categorizes the failure types.
Table 18: Failure case analysis for AgentTrace
Failure Type Count Percentage
Multiple plausible root causes 12 42.9%
Bug at unusual position 8 28.6%
Atypical causal structure 5 17.9%
Feature extraction issues 3 10.7%
Total 28 100%
The dominant failure mode is multiple plausible root causes: in scenario res_3867e9, both Step 2 (incorrect search query) and Step 3 (overly strict filter) contributed to the error; AgentTrace selected Step 2 while ground truth was Step 3. The second mode is bugs at unusual positions: scenarios with late-stage bugs (step 7+) have feature distributions that differ from the majority, causing AgentTrace to over-prioritize earlier steps. Potential improvements include position-adaptive weights, semantic similarity between error descriptions and node content, and multi-root-cause detection.
Appendix J Reproducibility
All code, data, and evaluation scripts are available at https://github.com/zwang000/AgentTrace, including benchmark generation scripts, 550 scenario JSON files with ground truth, all baseline implementations, and statistical analysis scripts. Table 19 lists all hyperparameters. Random seeds: benchmark generation (42), bootstrap resampling (12345), train/validation split (2024).

### 27 Mar 2026 v2

@@ 1 Introduction
While our benchmark scenarios are synthetically constructed, they are designed to capture common failure patterns observed in deployed multi-agent systems, such as early planning errors cascading through execution. Our evaluation across 550 failure scenarios in 10 domains shows that AgentTrace achieves high localization accuracy with sub-second latency, significantly outperforming both heuristic and LLM-based baselines. Figure 1 provides an overview of the framework.
Figure 1: Overview of the AgentTrace framework. (a) Execution logs from a multi-agent workflow are collected. (b) A causal graph is constructed with sequential, communication, and data dependency edges. (c) Backward BFS from the error node identifies candidate root causes. (d) Candidates are ranked using a weighted combination of five interpretable feature groups, with position features dominating ($w_{p}{=}0.70$).
@@ 3.4 Node Ranking Algorithm
$F_{i}(v)=\frac{1}{|G_{i}|}\sum_{j\in G_{i}}\hat{f}_{j}(v),\quad\hat{f}_{j}(v)=\frac{f_{j}(v)-\min_{u}f_{j}(u)}{\max_{u}f_{j}(u)-\min_{u}f_{j}(u)+\epsilon}$ (2)
where $\epsilon=10^{-8}$ prevents division by zero. The weights $w_{i}$ are determined via grid search on a held-out validation set of 50 scenarios (see Appendix B). Table 1 lists all 17 features across the five groups.
Table 1: Complete feature list with computation methods
Group ($w_{i}$) Feature Formula Range
Position (0.70) Normalized Position $\text{pos}(v)/|V|$ $[0,1]$
Depth Ratio $\text{depth}(v)/\max_{u}\text{depth}(u)$ $[0,1]$
Reverse Position $1-\text{pos}(v)/|V|$ $[0,1]$
Structure (0.20) Out-degree $|\{u:(v,u)\in E\}|/\max_{w}|\{u:(w,u)\in E\}|$ $[0,1]$
In-degree $|\{u:(u,v)\in E\}|/\max_{w}|\{u:(u,w)\in E\}|$ $[0,1]$
Reachability $|\text{descendants}(v)|/|V|$ $[0,1]$
Content (0.05) Error Keywords $\mathbb{1}[\text{``error''}\in\text{content}(v)]$ $\{0,1\}$
Uncertainty $\mathbb{1}[\text{``maybe''}\in\text{content}(v)]$ $\{0,1\}$
Length Anomaly $|\text{len}-\mu|/(3\sigma)$ (clipped) $[0,1]$
Keyword Density $\text{count}(\text{keywords})/\text{len}$ $[0,1]$
Flow (0.03) Agent Switch $\mathbb{1}[\text{agent}(v)\neq\text{agent}(\text{prev}(v))]$ $\{0,1\}$
Role Criticality $\text{role\_weight}(\text{agent}(v))$ $[0,1]$
Communication $\mathbb{1}[\text{type}(v)=\text{``message''}]$ $\{0,1\}$
Confidence (0.02) Stated Confidence $\text{conf}(v)$ if available, else 0.5 $[0,1]$
Hedging Score $\text{count}(\text{hedge\_words})/10$ (clipped) $[0,1]$
@@ 4.1 Scenario Generation
We created a benchmark of 550 synthetic multi-agent failure scenarios across 10 domains (Table 2). Each scenario includes a complete execution trace with 8–15 agent actions, a systematically injected bug at a specific node, and ground-truth annotation of the root cause node.
Table 2: Benchmark domains and scenario counts (550 total)
Domain N Domain N
Software Development 52 Healthcare Coordination 60
Customer Service 51 Legal Document Analysis 60
Research Analysis 51 Educational Tutoring 60
Planning & Scheduling 46 Financial Advisory 60
Financial Trading 50 DevOps Automation 60
4.2 Bug Injection and Ground Truth
We inject five categories of bugs: logic errors (incorrect conditionals, 30%), communication failures (message misinterpretation, 20%), data corruption (incorrect transformation, 20%), missing validation (skipped verification, 16%), and role confusion (acting outside expertise, 14%).
Bug injection follows a five-step process: (1) generate a correct execution trace from domain templates; (2) select a bug location (early steps 2–3: 60%, middle steps 4–6: 30%, late steps 7+: 10%); (3) assign a bug type; (4) inject erroneous content (e.g., flipped comparison operators, message truncation, swapped variable names); (5) propagate cascading effects to downstream steps. To prevent methods from exploiting explicit bug markers, we also produce a blind version with anonymized IDs and separated ground truth.
Each scenario’s ground truth was established through: (1) controlled bug injection at a known step, (2) verification that removing the bug prevents the error, and (3) confirmation that the bug’s effects propagate to the error node. The constructed causal graphs have on average 10.8 nodes ($\sigma{=}2.1$) and 14.2 edges ($\sigma{=}3.5$), with edges predominantly sequential (61%), followed by communication (27%) and data dependency (12%).
@@ 5.2 Baselines
  • LLM Analysis: GPT-4 with full trace and prompt engineering (details in Appendix E)
@@ 5.3 Main Results
Table 3 shows that AgentTrace significantly outperforms all baselines. McNemar’s test confirms significance ($p<10^{-10}$) for all comparisons, with Cohen’s h = 0.77 vs. LLM Analysis indicating a large practical difference.
Table 3: Root cause localization accuracy and statistical significance on the 550-scenario benchmark. 95% CIs from bootstrap resampling (B=10, [^0] B{=}10{,}000). All $p$-values $10^{-10}$ (McNemar’s test).
Method Hit@1 Hit@3 MRR $\chi^{2}$ Cohen’s h
Random 9.1% 27.3% 0.18 459.2 2.41
First Node 3.6% 10.9% 0.07 497.0 2.72
Last Node 12.7% 38.2% 0.25 436.5 2.29
LLM Analysis 68.5% 81.4% 0.74 102.5 0.77
AgentTrace 94.9% [92.9, 96.7] 98.4% [96.9, 99.4] 0.97
@@ 5.4 Ablation Study
Table 4 shows Hit@1 for all feature group combinations and the sensitivity to the position weight $w_{p}$. Position features alone reach 87.3%, while each additional group provides incremental gains. The optimal $w_{p}=0.70$ balances predictive power with discriminative information from other groups.
Table 4: Ablation study. Left: Hit@1 with different feature group combinations (P=Position, S=Structure, C=Content, F=Flow, E=Confidence). Right: Sensitivity of Hit@1 to position weight $w_{p}$.
Features Hit@1 $\Delta$ $w_{p}$ Hit@1
All (P+S+C+F+E) 94.9% 0.50 91.3%
P only 87.3% -7.6 0.60 93.5%
S only 34.5% -60.4 0.70 94.9%
C only 28.7% -66.2 0.80 93.8%
F only 15.2% -79.7 0.90 91.1%
E only 12.1% -82.8
P+S 92.4% -2.5
P+C 90.9% -4.0
P+F / P+E 89.5 / 88.7% -5.4 / -6.2
S+C / S+F 45.6 / 38.2% -49.3 / -56.7
P+S+C 93.8% -1.1
P+S+C+F 94.5% -0.4
5.5 Robustness Analysis
Table 5 stratifies results across three dimensions, demonstrating that AgentTrace is robust to bug type (93.6–95.8% Hit@1), trace length (93.9–95.5%), and bug position (92.7–96.1%).
Table 5: Robustness analysis: Hit@1 and MRR stratified by bug type, trace length, and bug position.
Stratification Category Hit@1 MRR
Bug Type Logic Error 95.8% 0.98
Communication Failure 93.6% 0.96
Data Corruption 94.5% 0.97
Missing Validation 94.3% 0.96
Role Confusion 94.8% 0.97
Trace Length 8–9 steps 95.5% 0.97
10–11 steps 95.0% 0.97
12–13 steps 94.2% 0.96
14–15 steps 93.9% 0.96
Bug Position Early (steps 2–3) 96.1% 0.98
Middle (steps 4–6) 93.3% 0.96
Late (steps 7+) 92.7% 0.95
5.6 Illustrative Examples
We present two representative scenarios to illustrate how AgentTrace operates.
Example 1: Software Development.
Task: Implement average calculation.
Step 1 [Planner]: Analyze requirements
  -> "Sum all numbers, count elements, divide"
Step 2 [Coder]: Write implementation
  -> "def average(nums): ..."
Step 3 [Coder]: Add edge case handling  [BUG]
  -> "if nums = []:  # BUG: = instead of =="
Step 4 [Reviewer]: Review code -> "Code looks correct."
Step 5 [Executor]: Run tests  [ERROR]
  -> "SyntaxError: invalid syntax at line 2"
Root cause: Step 3 (logic error). AgentTrace correctly identifies Step 3; the LLM baseline selects Step 5 (error node), illustrating its tendency to select the error manifestation rather than tracing back.
Example 2: Research Analysis.
Task: Analyze transformer architecture publications.
Step 1 [Searcher]: Query databases -> "150 papers"
Step 2 [Searcher]: Filter -> "25 relevant papers"
Step 3 [Analyzer]: Extract findings  [BUG]
  -> "Attention being replaced by MLPs"
     (actually: attention being enhanced)
Step 4 [Synthesizer]: "Field moving away from
     attention-based models..."
Step 5 [Writer]: "Transformers becoming obsolete"
Step 6 [Writer]: Finalize report  [ERROR]
  -> "Contradicts recent SOTA results"
Root cause: Step 3 (communication failure). AgentTrace correctly identifies Step 3; the LLM selects Step 4 (propagation step), demonstrating the challenge of distinguishing root causes from their downstream effects.
@@ 6 Discussion
Failure Case Analysis.
Among the 28 scenarios (5.1%) where AgentTrace failed to rank the correct root cause first, the dominant failure mode is multiple plausible root causes (12 cases, 42.9%): e.g., in scenario res_3867e9, both Step 2 (incorrect search query) and Step 3 (overly strict filter) contributed to the error. The second mode is bugs at unusual positions (8 cases, 28.6%): late-stage bugs (step 7+) have feature distributions that differ from the majority, causing AgentTrace to over-prioritize earlier steps. The remaining failures stem from atypical causal structures (5 cases) and feature extraction issues (3 cases). Potential improvements include position-adaptive weights and multi-root-cause detection.
@@ 7 Conclusion
Our benchmark generation code, evaluation scripts, and all experimental results are available at https://github.com/GeoffreyWang1117/AgentTrace/tree/iclr2026-aiwild-camera-ready. The benchmark includes complete execution traces, ground truth annotations, and baseline implementations.
@@ Appendix A Benchmark Generation Details
Appendix B Feature Weight Learning
@@ Appendix C Statistical Analysis Details
Appendix D Detailed Per-Domain Results
@@
Appendix E LLM Baseline Details
We use GPT-4 (gpt-4-0613) with temperature 0.0, max tokens 1024, top-pp 1.0, and zero frequency/presence penalties. The prompt instructs the model to analyze the full execution trace, identify causal relationships, and output only the step number of the root cause:
@@ 4. Consider: logic errors, miscommunication,
Error Type Count %
Selected error node 82 47.4
Off-by-one (adjacent) 45 26.0
Intermediate step 31 17.9
Completely incorrect 15 8.7
Total 173 100
Appendix F Runtime Performance Details
Tables 10 and 10 break down runtime by component and show scaling with trace length. Runtime scales approximately linearly: $T\approx 12.5n+5$ ms, where $n$ is the number of steps. All experiments were conducted on an AMD Ryzen 9 5950X (16-core, 3.4 GHz), 128 GB DDR4 RAM, NVMe SSD, Ubuntu 24.04 LTS, Python 3.13.
@@ Appendix F Runtime Performance Details
Appendix G Reproducibility
All code, data, and evaluation scripts are available at https://github.com/GeoffreyWang1117/AgentTrace/tree/iclr2026-aiwild-camera-ready, including benchmark generation scripts, 550 scenario JSON files with ground truth, all baseline implementations, and statistical analysis scripts. Table 11 lists all hyperparameters. Random seeds: benchmark generation (42), bootstrap resampling (12345), train/validation split (2024).